40x+3y+9z=35
15x+16y+11z=50.5
11x+10y+300z=370
Please help me to solve this, thank you!

Answer by Alan3354(69443)

40x+3y+9z=35
15x+16y+11z=50.5
11x+10y+300z=370
---------------------
Inconvenient coefficiencts.
---
I would use determinants, or Cramer's rule.
====
For the Det:
|40    3    9|
|15   16   11|
|11   10  300|

Det = 40*(16*300 - 10*11) - 3*(15*300 - 11*11) + 9*(15*10 - 11*16)
Det = 187600 - 13137 - 234
Det = 174229
=======================
Do you know how to do the rest?
You can check your answers using an online matrix solver.